Solve: z(p - q) = z² + (x + y²).
1
Expert's answer
2020-03-19T06:38:16-0400

"pz-qz=z^2+(x+y)^2\\
\frac{dx}{z}=\frac{dy}{-z}=\frac{dz}{z^2+(x+y)^2}\\
1) \frac{dx}{z}=\frac{dy}{-z}\implies\\
-dx=dy\implies -\int dx=\int dy\implies\\
-x+c_1=y \implies c_1=x+y\\
2) \frac{dx}{z}=\frac{dz}{z^2+(x+y)^2}
\implies\\
\frac{dx}{z}=\frac{dz}{z^2+c_1^2}\implies\\
dx=\frac{zdz}{z^2+c_1^2} \implies\\
\int dx=\int\frac{zdz}{z^2+c_1^2}\implies\\
\int dx=\frac{1}{2}\int\frac{d(z^2+c_1^2)}{z^2+c_1^2}\implies\\
\frac{1}{2}\ln|z^2+c_1^2|=x+\frac{1}{2}\ln|c_2|\implies\\
\ln|z^2+c_1^2|=2x+\ln|c_2|\implies\\
z^2+c_1^2=c_2e^{2x}\implies\\
z^2+(x+y)^2=c_2e^{2x}\implies\\
c_2=\frac{z^2+(x+y)^2}{e^{2x}}\\"

Solution of equation is

"F(x+y, \frac{z^2+(x+y)^2}{e^{2x}})=0"